Tim rolls a marble off a horizontal lab table with an initial speed of 2.5 m/s. The marble falls a vertical distance of 1.2 m to the floor below. Air resistance can be ignored.
a) How much time does the marble spend in the air?

b) What is the speed of the marble just before it strikes the floor?
1
Expert's answer
2020-12-10T11:06:23-0500

a)


"h=0.5gt^2\\1.2=0.5(9.8)t^2\\t=0.495\ s"

b)


"v=\sqrt{v_x^2+(gt)^2}\\v=\sqrt{2.5^2+(9.8\cdot0.495)^2}=5.46\frac{m}{s}"